OpenStudy (anonymous):

7th Grade Help ? ASAP.

OpenStudy (anonymous):

During a period of six months, Debbie's Dog-Walking business had monthly profits of $256, $318 , and $194. Two months had losses of $245 and $145. The net profit was $104. A. Write an equation to find the profit or loss for the month that was not included in the problem data. let n represent the n unknown profit or loss. B. Solve the equation you wrote in Part A. Show All Your Work. C. Was there a profit or loss for that month and for how much? Answer in a complete sentence

OpenStudy (jack1):

profits +256 +318 +194 minus losses -245 -145 plus unknown + u = +104 so total eqn = 256 +318 +194 -245 -145 +u = 104

OpenStudy (jack1):

can you add everything on the LHS of the equals sign to give one number?

OpenStudy (anonymous):

LHS?

OpenStudy (nathan917):

The left-hand side.

OpenStudy (anonymous):

I'm so lost, sorry I don't pay any attention in math class.

OpenStudy (nathan917):

That why we are here.

OpenStudy (nathan917):

@Jack1 You've got you smart score to 60 and on the way only 20 fans? :P

OpenStudy (nathan917):

your*

OpenStudy (nathan917):

Wow you just lost 2 fans how?

OpenStudy (jack1):

is that good? maybe i've just got an abrasive personality... and it comes across... in maths forums...?

OpenStudy (anonymous):

^ i thought y'all where helping not talking

OpenStudy (jack1):

we will, just first add every number together on the Left hand side of the equals sign

OpenStudy (anonymous):

you mean 256, 318, 194, 245, 195 and 104?

OpenStudy (jack1):

not 104, thats on the right of the equals sign

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

1208 ?

OpenStudy (jack1):

256 +318 +194 -245 -145 +u = 104 378 + u = 104 378 + u - 378 = 104 -378 u = 104 - 378

OpenStudy (jack1):

i see what you did there, but as the last 2 numbers were negative, you either need to add it as a negative number, or just simply subtract them from the others

OpenStudy (anonymous):

OH, sorry I didn't see they where negative.

OpenStudy (jack1):

all good, so whats your final answer for "u"?

OpenStudy (anonymous):

(/.\) I'm no good at this

OpenStudy (jack1):

ok so, u = 104 - 378 so it will be a negative number, which means that that month was also a loss so what's the answer...?

OpenStudy (jack1):

104 - 378 _______ =answer

OpenStudy (anonymous):

-274 ?

OpenStudy (jack1):

spot on

OpenStudy (jack1):

so you've go A you've done B and if you write it a a sentence, you've done C props, slaters

OpenStudy (anonymous):

lol yeah, actually. I'm more confused then ever but I'll figure what just happened sooner or later.

OpenStudy (jack1):

ok, all you did was make a profit loss equation we collected all of the data added together all of the months that we had profits subtracted all of the months that had losses and then called the month where we didn't know how much we made : U

OpenStudy (jack1):

and put all that on the LHS of the equals sign then on the RHS of the equals sign, we put how much money was made overall ($104) then reducing the equation, we solved for the unknown month

OpenStudy (jack1):

does that help a little...?

OpenStudy (anonymous):

yes thank you !

OpenStudy (jack1):

welcomes